# Disentangling Rich Dynamics from Feature Learning: A Framework for Independent Measurements

### v2

@@ Abstract
In machine learning, it is widely believed that dynamic feature transformation (the rich regime) enhances predictive performance. However, this link does not always hold, and existing richness measures rely on correlated factors — such as performance or parameter norms — which can complicate the analysis of feature learning. We introduce (1) a measure that quantifies the rich regime independently of performance, and (2) interpretable feature metrics for visualization. Leveraging low-rank bias, our approach generalizes neural collapse metrics and captures lazy-to-rich transitions (e.g., grokking) without relying on performance as a proxy. We reveal how batch normalization and training set size influence lazy/rich dynamics for VGG16 and ResNet18 on CIFAR-10/100, opening avenues for better understanding feature learning.
@@ 1 Introduction
In machine learning, feature learning is often viewed through two complementary perspectives: improvement of representations and non-linear training dynamics. The representation perspective emphasizes feature quality — how well it supports downstream tasks like classification and promotes generalization [1]. The dynamics perspective — often called rich regime in rich versus lazy training [2] — concerns the dynamic transformation of features beyond linear models. While dynamical richness frequently correlates with representation usefulness, rich dynamics reflect a preference (inductive bias) toward certain solutions, without necessarily benefiting all tasks [3]. For example, dynamic feature learning can impair the performance on an image classification task (Fig. 1).
The representation perspective of feature learning — central to deep learning’s success — remains poorly understood. However, its dynamic aspects, often shaped in practice by initialization and optimization, are well understood in the context of layerwise linear models (e.g., linear networks). Greedy dynamics with low-rank biases [4, 5] explain neural collapse [6] in vision tasks; delayed saturation dynamics [7] have been linked to emergence [8] and scaling laws [9] in large language models; and initialization-dependent dynamics [10, 11, 12, 13] offer insights into grokking [14] in arithmetic tasks (see Nam et al. [13] for a comprehensive overview).
To better understand the link between rich dynamics and representation improvement, we need independent metrics for dynamics and representations. However, existing dynamics metrics, often designed for purposes beyond practical richness measurement, can conflate learning dynamics with factors such as representation quality, labels, or the initial kernel.
@@
$\mathcal{D}_{LR}$downarrow $8.7\cdot 10^{-3}$ $6.3\cdot 10^{-1}$
@@
Figure 1: Rich dynamics $\neq$ better generalization. We trained a 4-layer MLP on label-encoded MNIST. (a): The first 10 pixels are encoded with true labels in training and random labels in testing; both the encoding and the image serve as valid features for training. (b): A full backpropagation model (rich) biases toward the encodings and generalizes poorly, while a last-layer-only-trained model (lazy) relies on the full image and generalizes better. Our low-rank-based metric $\mathcal{D}_{LR}$ (Eq. 4) quantifies the dynamical richness ($\mathcal{D}_{LR}\in[0,1]$ where 00 is richest). (c): Visualization framework (Eq. 5): (i) Cumulative contribution of last-layer features in spanning the target function $\Pi^{*}(k)$ — top 10 features are irrelevant in the rich model. (ii) Contribution to the learned function $\hat{\Pi}(k)$ — the rich model uses only the top 10 features, while the lazy model uses all. (iii) Relative feature norms $\rho_{k}/\rho_{1}$ — rich model concentrates on the top 10; lazy model decays more gradually. Test accuracies and $\mathcal{D}_{LR}$ values are shown in parentheses and square brackets, respectively. See Section 3.2 for details, and Appendix C for background and motivation.
In this paper, we propose a principled metric for dynamical richness that (1) generalizes neural collapse, (2) requires no access to labels or performance, and (3) operates in function space. We further introduce visualization methods that track feature quality, usage, and intensity throughout training. These tools reveal lazy-to-rich transitions (e.g., target downscaling [2, 3] and grokking) without referencing performance and help uncover new patterns, such as the effects of batch normalization and dataset size on rich dynamics. Together, they offer practical tools and conceptual clarity toward understanding feature learning.
@@
  • We propose a metric for dynamical richness independent of representational quality (Eq. 4) and formally show it recovers neural collapse as a special case (Propositions 1 and 2).
  • We introduce a visualization framework that disentangles feature quality and dynamical richness (Section 3.2), offering interpretable diagnostics for analyzing training dynamics.
  • Using this framework, we reveal known lazy-to-rich transitions without referencing performance. For example, it separately visualizes the representation improvement and lazy-to-rich transition in grokking (Fig. 4) and target downscaling (Fig. 2).
  • We reveal new empirical patterns, including a correlation between batch normalization and dynamical richness in VGG16 trained on CIFAR-100 and a correlation between feature quality and intensity in ResNet18 on CIFAR-100.
@@ 2 Set up and Background
For analytical tractability, we use MSE loss and supervised settings — e.g., balanced $C$-way classification or regression with scalar output — where target function entries are orthogonal and isotropic.
@@
For the input space $\mathcal{X}$, we train on $n$ samples from $q$, the underlying true distribution on the input space $\mathcal{X}$. The width of the last layer (or the linear model) is $p$. The dimension of the output function (and the output of neural network) is $C$. Denote $\hat{f}$ and $f^{*}$ as the learned function (at any point in training) and the target function, respectively. We use bra-ket notation $\braket{f}{g}:=\mathbf{E}_{x\sim q}[f(x)g(x)]$ to mean the inner product over the input distribution (Appendix B). See Appendix A for a glossary.
@@
Full experimental details and statistical significance for the tables are provided in Appendix F. All error bars represent one standard deviation.
@@
We define the feature map $\Phi:\mathcal{X}\rightarrow\mathbb{R}^{p}$ as the map from the input to the post-activations of the penultimate layer. For $1\leq k\leq p$, denote the $k^{\mathrm{th}}$ feature as the $k^{\mathrm{th}}$ entry $\Phi_{k}(x)$ of $\Phi(x)=\left[\Phi_{1}(x),\ldots,\Phi_{p}(x)\right]$. We define the feature kernel operator $\mathcal{T}:L^{2}(\mathcal{X})\rightarrow L^{2}(\mathcal{X})$, generalization of self-correlation matrix in function space, as
$\mathcal{T}[f](x^{\prime}):=\mathbf{E}_{x\sim q}\left[\sum_{k=1}^{p}\Phi_{k}(x)\Phi_{k}(x^{\prime})f(x)\right],\quad\quad\mathcal{T}=\sum_{k=1}^{p}\ket{\Phi_{k}}\bra{\Phi_{k}},$ (1)
where $L^{2}(\mathcal{X})$ is the Hilbert space of square-integrable functions on $\mathcal{X}$ (with distribution $q$). The second expression is $\mathcal{T}$ in bra-ket notation. Using Mercer’s theorem [15], we can decompose the operator into eigenfunctions and eigenvalues
$\mathcal{T}:=\sum_{k=1}^{p}\rho_{k}\ket{e_{k}}\bra{e_{k}},\quad\quad\mathcal{T}[e_{k}]=\rho_{k}e_{k},\quad\quad\braket{e_{k}}{e_{l}}=\delta_{kl},$ (2)
where $\rho_{k}\in\mathbb{R}$ is a non-negative eigenvalue and $e_{k}:\mathcal{X}\rightarrow\mathbb{R}$ is the orthonormal eigenfunction. The operator $\mathcal{T}$ and the eigenfunctions $e_{k}$ play a key role in kernel regression with feature map $\Phi$ (i.e. $K(x,x^{\prime})=\Phi(x)^{T}\Phi(x^{\prime})$). The linear model (kernel) has inductive bias toward larger (eigenvalue) eigenfunctions [16, 17, 18, 19, 20, 21] and generalizes better when larger eigenfunctions (larger $\rho_{k}$) also better describe the target function (larger $\braket{e_{k}}{f^{*}}$) — also known as the task-model alignment [16]. See Appendix C for a gentle introduction.
@@
Chizat et al. [2] introduced rich dynamics as deviations from the exponential saturation observed in linear models. Subsequent works [11, 22, 13] identified these dynamics as sigmoidal saturation or amplifying dynamics — arising from gradient descent in layerwise models — and connected them to phenomena like neural collapse [5], feature emergence [7], and grokking [11]. See Nam et al. [13] for a comprehensive overview.
@@
Low-rank hidden representations naturally emerge in the rich regime, as shown by studies across linear networks [4, 23, 24, 25, 26, 27], unconstrained feature models [5, 28], and matrix factorization [29, 30]. These works find that gradient dynamics decouple into a minimal number of modes, governed by the rank of the input-output correlation matrix (often that of the output $C$), causing gradients to concentrate on those modes and produce low-rank representations. In practice, the low-rank bias is observed as neural collapse [6], which consists of four conditions NC1-4 stated in Appendix D. Although often associated with improved representations, neural collapse does not consistently imply better generalization [31, 32, 33, 34], remaining largely a dynamical phenomenon [13]. See Kothapalli et al. [35] for an overview.
3 Richness measure and visualization method
We introduce a framework that consists of a dynamical richness measure and a visualization method. We will exploit the low-rank bias of rich dynamics to measure richness and show that it reduces to neural collapse as a special case. For visualization methods, we will generalize the natural metrics from kernel regression (linear models) to quantify the features’ cumulative quality $\Pi^{*}$, utilization $\hat{\Pi}$, and relative intensity $\rho_{k}/\rho_{1}$ (Eq. 5), providing objective metrics for the evolution of features.
@@ 3.1 Low rank bias as richness measure
We define the learned function space ^$\hat{\mathcal{H}}=span\{\hat{f}_{1},\ldots,\hat{f}_{C}\}$ where $\hat{f}_{k}:\mathcal{X}\rightarrow\mathbb{R}$ is the $k^{\mathrm{th}}$ entry of the learned function (at any given time). In an ideal rich dynamics scenario, only the minimal number of features are learned throughout the dynamics and are sufficient to express (linearly span) the learned function space. Leveraging this idea, we define the minimal projection operator $\mathcal{T}_{MP}$.
@@
$\mathcal{T}_{MP}[u]=a_{1}\braket{\mathbf{1}}{u}\mathbf{1}+a_{2}P_{\hat{\mathcal{H}}}(u)\quad\text{for all }u\in L^{2}(\mathcal{X}),$ (3)
where $a_{1},a_{2}>0$, $\mathbf{1}$ is a constant function, and $P_{\hat{\mathcal{H}}}$ is the orthogonal projection onto $\hat{\mathcal{H}}$.
Ignoring the constant function (setting $a_{1}=0$), whose discussion is deferred to Appendix D, the $\mathcal{T}_{MP}$ is (up to a constant scale) a projection operator $P_{\hat{\mathcal{H}}}$ that removes all components orthogonal to $\hat{\mathcal{H}}$. If $\mathcal{T}$ is $\mathcal{T}_{MP}$, the last-layer features span only a $C$-dimensional space that matches the learned function space, reflecting the low-rank structure characteristic of rich dynamics. We thus define the low rank measure $\mathcal{D}_{LR}$ as the similarity between the $\mathcal{T}$ — defined by the current features — and the MP-operator $\mathcal{T}_{MP}$ — defined by the current learned function — as a metric for richness:
$\mathcal{D}_{LR}:=1-CKA(\mathcal{T},\mathcal{T}_{MP}),$ (4)
where $CKA$ is the centered kernel alignment [36] with bounded value in $[0,1]$. Because $CKA$ is normalized and uses centered (zero-mean) alignment, the metric remains consistent for any $a_{1}$ and $a_{2}$ (Section B.3). We subtract the $CKA$ measure from $1$ so that lower values indicate richer dynamics, consistent with more widely used metrics of richness. The novelty of our metric $\mathcal{D}_{LR}$ lies in using the minimum projection operator $\mathcal{T}_{MP}$.
@@
Our metric (Eq. 4) estimates how many independent feature directions exist beyond those used to express the learned function. In rich dynamics, only the minimal necessary features are learned and used ($\mathcal{D}_{LR}=0$); the earlier layers are sufficiently feature-learned to express the learned function. An excess of features suggests a bottleneck —such as limited expressivity in earlier layers— hindering feature learning of earlier layers.
Connection to neural collapse.
Suppose that the empirical distribution coincides with the true distribution and the neural network perfectly classifies all labels with one-hot vectors. We show that if $\mathcal{T}$ is an MP-operator, then the NC1 and NC2 conditions (Appendix D) of neural collapse hold. Let dagger $\dagger$ denote pseudo inverse, $A_{i}$ the set of datapoints with class label $i$, so the $i^{\mathrm{th}}$ learned function $\hat{f}_{i}$ coincides with the one-hot indicator function for $A_{i}$. Following Papayan et al. [6], we define the $i^{\mathrm{th}}$ class mean vector as $\mu_{i}:=\mathbf{E}_{x\in A_{i}}[\Phi(x)]$ and the global mean vector as $\bar{\mu}:=C^{-1}\sum_{i=1}^{C}\mu_{i}$.

### v3

# Decoupling Dynamical Richness from Representation Learning: Towards Practical Measurement

@@ Abstract
Dynamic feature transformation (the rich regime) does not always align with predictive performance (better representation), yet accuracy is often used as a proxy for richness, limiting analysis of their relationship. We propose a computationally efficient, performance-independent metric of richness grounded in the low-rank bias of rich dynamics, which recovers neural collapse as a special case. The metric is empirically more stable than existing alternatives and captures known lazy-to-rich transitions (e.g., grokking) without relying on accuracy. We further use it to examine how training factors (e.g., learning rate) relate to richness, confirming recognized assumptions and highlighting new observations (e.g., batch normalization promotes rich dynamics). An eigendecomposition-based visualization is also introduced to support interpretability, together providing a diagnostic tool for studying the relationship between training factors, dynamics, and representations.
@@ 1 Introduction
In machine learning, feature learning is often viewed through two complementary perspectives: improvement of representations and non-linear training dynamics. The representation perspective emphasizes feature quality — how well it supports downstream tasks like classification and promotes generalization (Bengio et al., 2013). The dynamics perspective — often called rich regime in rich versus lazy training (Chizat et al., 2019) — concerns the dynamic transformation of features beyond linear models. While dynamical richness frequently correlates with representation usefulness, rich dynamics reflect a preference (inductive bias) toward certain solutions, without necessarily benefiting all tasks (Geiger et al., 2020; Göring et al., 2025). For example, dynamic feature learning can impair the performance on an image classification task (Fig.˜1).
The representation perspective of feature learning — central to deep learning’s success — remains poorly understood. However, its dynamic aspects, often shaped in practice by initialization and optimization, are well understood in the context of layerwise linear models (e.g., linear networks). Greedy dynamics with low-rank biases (Saxe et al., 2014; Mixon et al., 2020) explains neural collapse (Papyan et al., 2020) in vision tasks; delayed saturation dynamics (Nam et al., 2024) has been linked to emergence (Brown et al., 2020) and scaling laws (Kaplan et al., 2020) in large language models; and initialization-dependent dynamics (Kumar et al., 2024; Kunin et al., 2024; Lyu et al., 2024) offers insights into grokking (Power et al., 2022) in arithmetic tasks. See  Nam et al. (2025) for a comprehensive overview.
To better understand the link between rich dynamics and representation improvement, we need independent metrics for dynamics and representations. However, prior dynamics metrics, while insightful, are not optimized for practical richness measurement (Section˜4.1). In this paper, we propose a dynamical richness metric that compares the activations before and after the last layer. The metric (1) generalizes neural collapse, (2) is computationally cheap, (3) is performance-independent, and (4) operates in the function space. We demonstrate that our metric empirically tracks dynamical richness without referencing performance, and assess its robustness against existing metrics. Finally, we experiment our metric, with complementary visualization, across various setups, highlighting its potential as a tool for uncovering new empirical insights.
@@
$\mathcal{D}_{LR}$ $\downarrow$ 0.0087 0.63
@@
Figure 1: Rich dynamics $\neq$ better generalization. We trained a 4-layer MLP on label-encoded MNIST. (a): The first 10 pixels are encoded with true labels in training and random labels in testing; both the encoding and the image serve as valid features for training. (b): A full backpropagation model (rich) biases toward the encodings and generalizes poorly, while a last-layer-only-trained model (lazy) relies on the full image and generalizes better. Our low-rank-based metric $\mathcal{D}_{LR}$ (Eq.˜4) quantifies the dynamical richness ($\mathcal{D}_{LR}\in[0,1]$ where 0 is richest) independent of the performance. (c): Complementary visualization method (Eq.˜5): (i) Cumulative contribution of last-layer features in expressing the target function — top 10 features are irrelevant in the rich model. (ii) Contribution to the learned function — the rich model uses only the top 10 features, while the lazy model uses all. (iii) Relative feature norms — rich model concentrates on the top 10; lazy model decays more gradually. Test accuracies and $\mathcal{D}_{LR}$ values are shown in parentheses and square brackets, respectively. See Section˜5.1 for details, and Appendix˜C for background and motivation.
@@
  • We introduce the Minimum projection operator $\mathcal{T}_{MP}$ (Definition˜1) and define the dynamical low-rank measure $\mathcal{D}_{LR}$ (Eq.˜4), a lightweight and performance-independent metric.
  • We show that $\mathcal{D}_{LR}$ reduces to neural collapse as a special case, connecting our formulation to a well-studied phenomenon while extending it to settings without labels.
  • We empirically confirm that $\mathcal{D}_{LR}$ captures known lazy-to-rich transitions such as grokking and target downscaling (Table˜2), aligning with theoretical expectations while giving a direct evaluation without referencing performance. We also empirically assess its robustness against alternatives.
  • We demonstrate the utility of our metric by showing how various training factors (e.g., architecture and learning rate) relate to richness and performance (Table˜3). Our metric explicitly confirms prior assumptions and uncovers new observations, such as batch normalization shifting VGG-16 on CIFAR-100 from lazy to rich dynamics. We further adapt eigendecomposition to visualize the metric for improved interpretability.
@@ 2 Setup and Background
For analytical tractability, we use MSE loss and supervised settings where target function entries are orthogonal and isotropic (e.g., balanced $C$-way classification or regression with scalar output).
@@
For the input space $\mathcal{X}$, we train on $n$ samples from $q$, the underlying true distribution on the input space $\mathcal{X}$. The width of the last layer (or the linear model) is $p$. The dimension of the output function (and the output of the neural network) is $C$. Denote $\hat{f}$ and $f^{*}$ as the learned function (at any point in training) and the target function, respectively. We use bra-ket notation $\braket{f|g}:=\mathbf{E}_{x\sim q}[f(x)g(x)]$ to mean the inner product over the input distribution (Appendix˜B). See Appendix˜A for a glossary.
@@
Full experimental details, link to the source code, and statistical significance for the tables are provided in Appendix˜H. All error bars represent one standard deviation.
@@
We define the feature map $\Phi:\mathcal{X}\rightarrow\mathbb{R}^{p}$ as the map from the input to the post-activations of the penultimate layer. For $1\leq k\leq p$, denote the $k^{\mathrm{th}}$ feature as the $k^{\mathrm{th}}$ entry $\Phi_{k}(x)$ of $\Phi(x)=\left[\Phi_{1}(x),\ldots,\Phi_{p}(x)\right]$. We define the feature kernel operator $\mathcal{T}:L^{2}(\mathcal{X})\rightarrow L^{2}(\mathcal{X})$, generalization of self-correlation matrix in the function space, as
$|,\mathcal{T}[f](x^{\prime}):=\mathbf{E}_{x\sim q}\left[\sum_{k=1}^{p}\Phi_{k}(x)\Phi_{k}(x^{\prime})f(x)\right],\quad\quad\mathcal{T}=\sum_{k=1}^{p}\ket{\Phi_{k}}\bra{\Phi_{k}}$, (1)
where $L^{2}(\mathcal{X})$ is the Hilbert space of square-integrable functions on $\mathcal{X}$ (with distribution $q$). The second expression is $\mathcal{T}$ in bra-ket notation (Appendix˜B). Using Mercer’s theorem (Mercer, 1909), we can decompose the operator into eigenfunctions and eigenvalues
[…]
where $\rho_{k}\in\mathbb{R}$ is a non-negative eigenvalue and $e_{k}:\mathcal{X}\rightarrow\mathbb{R}$ is the orthonormal eigenfunction. The operator $\mathcal{T}$ and the eigenfunctions $e_{k}$ play a key role in kernel regression with feature map $\Phi$ (i.e. $K(x,x^{\prime})=\Phi(x)^{T}\Phi(x^{\prime})$). The linear model (kernel) has inductive bias toward eigenfunctions corresponding to larger eigenvalues (Bordelon et al., 2020; Jacot et al., 2020; Spigler et al., 2020; Canatar et al., 2021; Cui et al., 2021; El Harzli et al., 2024), and generalizes better when eigenfunctions with larger corresponding eigenvalues also better describe the target function (larger $\braket{e_{k}|f^{*}}$) — also known as the task-model alignment (Bordelon et al., 2020). See Appendix˜C for a gentle introduction.
@@
Chizat et al. (2019) introduced rich dynamics as deviations from the exponential saturation observed in linear models. Subsequent works (Kunin et al., 2024; Dominé et al., 2025; Nam et al., 2025) identified these dynamics as sigmoidal saturation or amplifying dynamics — arising from gradient descent in layerwise models — and connected them to phenomena like neural collapse (Mixon et al., 2020), feature emergence (Nam et al., 2024), and grokking (Kunin et al., 2024). See Nam et al. (2025) for a comprehensive overview.
@@
Low-rank hidden representations naturally emerge in the rich regime, as shown by studies across linear networks (Saxe et al., 2014; Ji and Telgarsky, 2019; Arora et al., 2019a; Lampinen and Ganguli, 2019; Gidel et al., 2019; Tarmoun et al., 2021), unconstrained feature models (Mixon et al., 2020; Fang et al., 2021), and matrix factorization (Arora et al., 2019b; Li et al., 2020). These works find that gradient dynamics decouple into a minimal number of modes, governed by the rank of the input-output correlation matrix (often that of the output $C$), causing gradients to concentrate on those modes and produce low-rank representations. In practice, the low-rank bias is observed as neural collapse (Papyan et al., 2020), which consists of four conditions NC1-4 stated in Appendix˜D. Although often associated with improved representations, neural collapse does not consistently imply better generalization (Zhu et al., 2021; Nguyen et al., 2022; Hui et al., 2022; Su et al., 2023), remaining largely a dynamical phenomenon. See Kothapalli et al. (2022) for an overview.
Change in NTK as measure of rich dynamics.
Beyond performance, changes in the Neural tangent kernel (NTK, Jacot et al. (2018)) before and after training are among the most commonly used proxies for rich dynamics (Chizat et al., 2019; Atanasov et al., 2021; Kumar et al., 2024). Although theoretically well-founded, NTK-based measures scale with the total number of parameters and are computationally infeasible even for moderate vision models. For this reason, we omit NTK from our analysis and we focus only on the feature kernel of the last layer (Eq.˜1).
3 The richness measure $\mathcal{D}_{LR}$ (Minimum Projection Alignment)
We introduce a dynamical richness measure that exploits the low-rank bias of rich dynamics and show that it reduces to neural collapse as a special case.
@@ 3.1 Low rank bias as richness measure
We define the learned function space $\hat{\mathcal{H}}=\mathrm{span}\{\hat{f}_{1},\ldots,\hat{f}_{C}\}$ where $\hat{f}_{k}:\mathcal{X}\rightarrow\mathbb{R}$ is the $k^{\mathrm{th}}$ entry of the network’s learned function (at any given time). In an ideal rich dynamics scenario, only the minimal number of features are learned throughout the dynamics and are sufficient to express (linearly span) the learned function space. Leveraging this idea, we define the minimal projection operator $\mathcal{T}_{MP}$.
@@
$\mathcal{T}_{MP}[u]=a_{1}\braket{\mathbf{1}|u}\mathbf{1}+a_{2}P_{\hat{\mathcal{H}}}(u)\quad\text{for all }u\in L^{2}(\mathcal{X}),$ (3)
where $a_{1},a_{2}>0$, $\mathbf{1}\in L^{2}(\mathcal{X})$ is a constant function, and $P_{\hat{\mathcal{H}}}:L^{2}(\mathcal{X})\rightarrow L^{2}(\mathcal{X})$ is the orthogonal projection in $L^{2}(\mathcal{X})$ onto $\hat{\mathcal{H}}$.
Ignoring the constant function (setting $a_{1}=0$), whose discussion is deferred to Appendix˜D, the $\mathcal{T}_{MP}$ is (up to a constant scale) a projection operator $P_{\hat{\mathcal{H}}}$ that removes all components orthogonal to $\hat{\mathcal{H}}$. If $\mathcal{T}$ is $\mathcal{T}_{MP}$, the last-layer features span only a $C$-dimensional space that matches the learned function space, reflecting the low-rank structure characteristic of rich dynamics. We thus define the low rank measure $\mathcal{D}_{LR}$ as the similarity between the $\mathcal{T}$ — defined by the current features — and the MP-operator $\mathcal{T}_{MP}$ — defined by the current learned function — as a metric for richness:
$\boxed{\mathcal{D}_{LR}:=1-CKA(\mathcal{T},\mathcal{T}_{MP}),}$ (4)
where $CKA$ is the centered kernel alignment (Kornblith et al., 2019) with bounded value in $[0,1]$. Because $CKA$ is normalized and uses centered (zero-mean) alignment, the metric remains consistent for any $a_{1}$ and $a_{2}$ (Section˜B.3). We subtract the $CKA$ measure from $1$ so that lower values indicate richer dynamics, consistent with more widely used metrics of richness. The novelty of our metric $\mathcal{D}_{LR}$ lies in defining the minimum projection operator $\mathcal{T}_{MP}$ and comparing it to $T$ to quantify dynamical richness, not in using well-established CKA. Notably, CKA is primarily used to compare the NTK before and after training (Chizat et al., 2019; Kumar et al., 2024; Baratin et al., 2021).
@@
Our metric (Eq.˜4) compares the activations before ($\mathcal{T}$, at most rank $p$) and after ($\mathcal{T}_{MP}$, rank $C$) the last layer. In rich dynamics, low-rank bias dictates that only the minimal necessary $C$ out of $p$ features are learned and used before the final layer ($\mathcal{D}_{LR}=0$). In this case, the final layer performs no additional processing because earlier layers have already completed the task. Conversely, an excess of features indicates a bottleneck — such as limited expressivity in earlier layers — that prevents the full manifestation of low-rank bias.
Time complexity.
The algorithm is highly efficient. For $n$ samples (from either train or test), it requires $n$ forward passes to record activations before and after the final layer, producing matrices of size $n\times p$ and $n\times C$, where $p$ is the last-layer width and $C$ the number of classes. $\mathcal{D}_{LR}$ is then computed in $\mathcal{O}(npC)$. Since typically $n\gg p\gg C$, setting $n\approx\mathcal{O}(p)$ suffices, reducing complexity to $\mathcal{O}(p^{2}C)$. In standard models with $p\approx 10^{3}$, this is far cheaper than NTK-based methods, which scale quadratically with the total number of parameters. See Appendix˜E for details.
3.2 Connection to neural collapse
Suppose that the empirical distribution coincides with the true distribution and the neural network perfectly classifies all labels with one-hot vectors. We show that if $\mathcal{T}$ is an MP-operator, then the NC1 and NC2 conditions (Appendix˜D) of neural collapse hold. Let dagger $\dagger$ denote pseudo inverse, $A_{i}$ the set of datapoints with class label $i$, so the $i^{\mathrm{th}}$ learned function $\hat{f}_{i}$ coincides with the one-hot indicator function for $A_{i}$. Following Papyan et al. (2020), we define the $i^{\mathrm{th}}$ class mean vector as $\mu_{i}:=\mathbf{E}_{x\in A_{i}}[\Phi(x)]$ and the global mean vector as $\bar{\mu}:=C^{-1}\sum_{i=1}^{C}\mu_{i}$.